OpenStudy (anonymous):

dynamical system B(x)=2x if 0≤x≤1/2 B(x)=2x−1 if 1/2≤x≤1 then find a formula for D^n (x) ????????????????

OpenStudy (anonymous):

\[B(x)=\begin{cases}2x&\text{if }0\le x\le\dfrac{1}{2}\\2x-1&\text{if }\dfrac{1}{2}\le x\le 1\end{cases}\] What do you mean by \(D^n(x)\) ? Is that the \(n\)-th derivative?